# Freeze-in and Freeze-out in Right-Handed Neutrino Extended MSSM with Seesaw

### v1

@@ IV.1 Numerical procedure
The model files were generated with Sarah v4.14.5 [26, 27] and the particle spectrum was generated with SPheno v4.0.5 [28, 29]. Relic abundance calculations were performed using micrOMEGAs v5.3.41[30]. All our calculations were carried out in Feynman gauge. It is important to mention that micrOMEGAs v 5.3.41 has errors in computing the correct freeze-in values in unitary gauge (via ForceUG option) and gives inconsistent results with increasing reheating temperature (TR option). We present two benchmark points, BP1 and BP2, where the sneutrino (our freeze-in candidate) appears as LSP with different masses. The relevant parameters are listed in Table 1.
Using BP1, we generated 800 points by randomly sampling the parameter range shown in Table 1. In total, 12 parameters were varied, including $M^{2}_{\nu,N}$ and $y^{\nu}_{i,j}$. These ranges were chosen such that the total relic density remains within 10% of the observed DM relic density. In addition, we ensured that the sum of the left-handed neutrino masses lies between $0.06~\mathrm{eV}\text{ and }0.09~\mathrm{eV}$ [31] for all benchmark points. Fig. 5 was generated using this dataset. Similarly, Fig. 6 was obtained from a dataset generated from a random scan of about 46,000 points without performing relic abundance calculations using micrOMEGAs v 5.3.41. We will discuss these Figures in detail in the next section.
@@ IV.2 Relic density
Figure 3: Relic density as a function of scanned parameters, $\mu$ and $M^{2}_{\nu,N}$ for BP1. The kinks indicate changes in the nature of the LSP from neutralino to sneutrino as we increase the parameter.
Figure 4: Relic density as a function of scanned parameters, $y^{\nu}_{1,1}$ and $y^{\nu}_{2,1}$, demonstrating a quadratic dependence across BP1.
@@
In the final scan, many processes involving neutrinos were excluded, as they require excessive computational time while contributing only marginally. In addition, we neglected any process with a contribution to the relic density below $10^{-6}$. Taken together, these contribute at most about $3\times 10^{-3}$ to the relic density, corresponding to a maximum effect of about 3%. For reference, we have taken the present DM relic abundance to be $0.1206\pm 0.0021$ [32].
@@ IV.3 Direct detection
We also examined the possibility of a higgsino-like dark matter candidate. However, we were unable to identify any viable benchmark point consistent with the relic density and existing constraints. Moreover, with the most recent bounds from direct-detection experiments [2], such scenarios are now entirely ruled out in the models discussed here. For the case where the sneutrino is LSP, the spin-independent nucleon cross-section is of the order $10^{-90}\text{cm}^{2}$, unobservable in direct detection.
As discussed in Sec. III, higgsino-like dark matter is excluded by direct detection unless the gauginos are very heavy. In that case, the lightest neutralino is an almost pure higgsino of the form $\tilde{\chi}^{0}_{1}\simeq\frac{1}{\sqrt{2}}(\tilde{H}^{0}_{u}-\tilde{H}^{0}_{d})$, for which the cancellation in Eq. 19 suppresses the direct-detection cross section. With the most recent bounds by the LZ experiment, the electroweak gauginos would need to have masses of at least $15$ TeV. Cases with very heavy gauginos have also been studied (see e.g. [33]), as they provide a benchmark for inelastic direct detection. However, such a heavy spectrum would lead to sizable loop-induced corrections to Eq. 1 if we assume the gluino to lie in the same mass range as the electroweakinos (see e.g. discussion in [7]). In addition, the loop corrections to the Higgs mass would tend to become too large.
@@ V Sterile neutrino constraints
Figure 5: Mixing angle–mass plane for a right-handed (sterile) neutrino: $\sin^{2}\!\bigl(2\theta\bigr)$ vs. $m_{N}\,(\mathrm{GeV})$. The model points are distributed based on the total relic abundance by freeze-in and freeze-out. The background colour map shows the sterile-neutrino relic abundance $\Omega_{N}h^{2}$ from the Dodelson–Widrow mechanism.
@@
There are several direct and indirect constraints on right-handed (sterile) neutrinos. Collider searches [34, 35] and studies of kinematical endpoints in nuclear decays typically require large mixing between active and sterile neutrinos. Such mixing cannot be realised within the type-I seesaw model and would require so large Yukawa couplings in the inverse and linear seesaw models that the sneutrino component would lead to overabundance in our scenario. In addition, short baseline oscillation experiments rule out very light sterile neutrinos below a few eV [36].
The indirect constraints are more severe. Light sterile neutrinos are relatively long-lived. They decay to active neutrinos via $N\rightarrow 3\nu$ if no other channels are kinematically available. The consequences of this decay depend on the sterile neutrino lifetime. If the lifetime is of the order of one second, the decays will produce active neutrinos carrying energies higher than those expected from the thermal distribution, which can have an impact on Big Bang nucleosynthesis (BBN) [37]. Successful BBN therefore requires the sterile neutrinos either to decay well before BBN, allowing neutrinos to thermalise, or to be sufficiently long-lived.
If the sterile neutrinos have a lifetime longer than the age of the Universe, they constitute a (decaying) part of the dark matter. The dominant mechanism for sterile neutrino dark matter generation is oscillations between active and sterile neutrinos known as the Dodelson–Widrow mechanism [19]. In this case, the relic density depends primarily on the sterile neutrino mass and the mixing angle between the active and sterile flavours. The relic density can be approximately given by [38, 39]:
@@
where $m_{N}$ is the sterile neutrino mass and $\theta$ the mixing angle between the active and sterile neutrinos. The relic density depends rather mildly on model parameters since increasing the sterile neutrino mass leads to a smaller mixing angle.
For all our benchmarks, we find that the lifetime of sterile neutrinos is of the order $10^{42}$ GeV-1 ($2.087\times 10^{10}$ yrs), which is more than the age of the universe ($1.38\times 10^{10}$ yrs). Thus, they are stable on cosmological scales and contribute to the DM abundance.
As discussed earlier, the production of the right amount of sneutrino dark matter through freeze-in requires the neutrino Yukawa couplings of the order $\mathcal{O}(10^{-12})$. Eq. 5 implies that $m_{N}=(y^{\nu}v\sin\beta)^{2}/2m_{\nu}$. The heaviest active neutrino must be at least $0.04$ eV which leads to sterile neutrino masses in the range 552020 eV, not much above the lowest allowed values by short baseline oscillation experiments. The associated Dirac mass term is of the order of one electronvolt which leads to an active-sterile mixing angle $\theta\gtrsim 10^{-2}$.
Linear and inverse seesaw models allow to suppress this mixing by letting the sterile neutrinos to be more massive. In the linear seesaw model, however, this comes with a price as this introduces additonal interactions between the sterile sneutrinos and SM particles, which makes the freeze-in contribution too large as discussed earlier.
In the inverse seesaw, the mass matrix has the form
@@
If we wish to have heavier sterile neutrinos while keeping the light neutrino mass constant, the factor $\mu_{S}/M_{N}$ must be greater than one. The more we increase the sterile neutrino mass, the larger $\mu_{S}$ must become. For proper suppression of the mixing $\mu_{S}/M_{N}\gtrsim 5$ would be required. When $\mu_{S}$ is significantly larger than $M_{N}$, the lower $2\times 2$ block of the inverse seesaw mass matrix becomes equivalent to that of the type-I seesaw. Then we get sterile neutrino masses $M_{N}^{2}/\mu_{S}$ for $N$ and $\mu_{S}$ for $S$, which means that we end up having the top $2\times 2$ block and type-I seesaw mechanism. Hence, the inverse seesaw mechanism too cannot successfully suppress the Dodelson–Widrow contribution to the relic density. The only model where there is no problem is that of Dirac neutrinos, since there are no right-handed neutrinos.
In Fig. 5, we plot the mass of the sterile neutrinos (with the largest mixing angle) against $\sin^{2}2\theta$ using points from the type-I seesaw model. The red dots correspond to points that yield the correct relic density with an error margin of $10\%$, accounting for numerical or package-related uncertainties. The black square and blue cross indicate benchmarks that lead to DM overabundance and underabundance, respectively. As the Figure shows, all the points predict a relic density greater than 3, $\Omega_{N}h^{2}>3$, denoted by the reference line.
Another feature visible in Fig. 5 is the parabolic void in the lower left region of the plot, due to the lower bound on the Yukawa coupling specified in Table 1. By reducing this lower limit by half, this parabolic void disappears. Nonetheless, such points are excluded since they predict excessive contribution to DM relic density from sterile neutrinos via Dodelson–Widrow mechanism.
The overabundance is mainly driven by the heaviest right-handed neutrino. For example, $m_{N}=50$ eV and $\sin^{2}2\theta\gtrsim 10^{-4}$ yields $\Omega h^{2}\gtrsim 2$. In the type-I seesaw, it is impossible to suppress this mixing while preserving a viable neutrino mass spectrum. Fig. 6 illustrates this through a numerical scan, where the Yukawa coupling values are now relaxed to $10^{-15}$ and the sterile neutrino masses down to $0.1~\mathrm{eV}$. Points that allow smaller dark matter contributions from the right-handed neutrinos are excluded by oscillation experiments (lower right region), while the allowed points lead to a clear overabundance (upper middle). The experiments included in the blue contour are Bugey-3[40], NEOS[41], DANSS[42], Daya Bay[43], PROSPECT-I[44] and KATRIN[45].
In summary, although the combination of freeze-out and freeze-in can yield an acceptable relic density, the sterile neutrino contribution is always too large. If sterile neutrinos were made sufficiently heavy with a sufficiently short lifetime, the couplings between the SM and sneutrinos would become too large to be viable.
@@ VI Conclusions
We showed that an acceptable relic density requires very small neutrino Yukawa couplings and a sterile neutrino mass scale so low that the sterile neutrinos become cosmologically stable. In the region of parameter space where the relic density from the superpartners is viable, the production of sterile neutrinos in the early Universe via the Dodelson–Widrow mechanism leads to overabundance. Therefore, a scenario with freeze-in dark matter from sterile sneutrinos at the electroweak scale does not work for type-I, inverse, or linear seesaw mechanisms. For Dirac neutrinos, however, this scenario remains possible, as there are no sterile neutrino mass eigenstates that would lead to overabundance.
@@ Acknowledgments
TG acknowledges Magnus Ehrnrooth foundation for funding his doctoral studies. MH acknowledges the support from the Research Council of Finland (grants #342777 and #371542). HW acknowledges the support from Carl Trygger Foundation (grant CTS18:164) and Ruth and Nils-Erik Stenbäck’s Foundation. The authors wish to thank the Finnish Computing Competence Infrastructure (FCCI) for supporting this project with computational and data storage resources.

### v2

@@ IV.1 Numerical procedure
The model files were generated with Sarah v4.14.5 [27, 28] and the particle spectrum was generated with SPheno v4.0.5 [29, 30]. The relic abundance calculations were performed using micrOMEGAs v5.3.41[31]. All our calculations were carried out in Feynman gauge. It is important to mention that micrOMEGAs v 5.3.41 has errors in computing the correct freeze-in values in unitary gauge (via ForceUG option) and gives inconsistent results with increasing reheating temperature (TR option). We present two benchmark points, BP1 and BP2, where the sneutrino (our freeze-in candidate) appears as the LSP with different masses. The relevant parameters are listed in Table 1.
Using BP1, we generated 800 points by randomly sampling the parameter range shown in Table 1. In total, 12 parameters were varied, including $M^{2}_{\nu,N}$ and $y^{\nu}_{i,j}$. These ranges were chosen such that the total relic density remained within 10% of the observed DM relic density. In addition, we ensured that the sum of the left-handed neutrino masses lie between $0.06~\mathrm{eV}\text{ and }0.09~\mathrm{eV}$ [32] for all benchmark points. Fig. 5 was generated using this dataset. Similarly, Fig. 6 was obtained from a dataset generated from a random scan of about 46,000 points without performing relic abundance calculations using micrOMEGAs v 5.3.41. We will discuss these figures in detail in the next section.
@@ IV.2 Relic density
Figure 3: Relic density as a function of the scanned parameters $\mu$ and $M^{2}_{\nu,N}$ for BP1. The kinks indicate changes in the nature of the LSP from neutralino to sneutrino as we increase the parameter.
Figure 4: Relic density as a function of the scanned parameters $y^{\nu}_{1,1}$ and $y^{\nu}_{2,1}$, demonstrating a quadratic dependence across BP1.
@@
In the final scan, many processes involving neutrinos were excluded, as they required excessive computational time while contributing only marginally. In addition, we neglected any process with a contribution to the relic density below $10^{-6}$. Taken together, these contribute at most about $3\times 10^{-3}$ to the relic density, corresponding to a maximum effect of about 3%. For reference, we have taken the present DM relic abundance to be $0.1206\pm 0.0021$ [33].
@@ IV.3 Direct detection
We also examined the possibility of a Higgsino-like dark matter candidate. However, we were unable to identify any viable benchmark point consistent with the relic density and existing constraints. Moreover, with the most recent bounds from direct-detection experiments [2], such scenarios are now entirely ruled out in the models discussed here. For the case where the sneutrino is an LSP, the spin-independent nucleon cross section is of the order $10^{-90}\text{cm}^{2}$, unobservable in direct-detection.
As discussed in Section III, Higgsino-like dark matter is excluded by direct-detection unless the gauginos are very heavy. In that case, the lightest neutralino is an almost pure Higgsino of the form $\tilde{\chi}^{0}_{1}\simeq\frac{1}{\sqrt{2}}(\tilde{H}^{0}_{u}-\tilde{H}^{0}_{d})$, for which the cancellation in Eq. 19 suppresses the direct-detection cross section. With the most recent bounds by the LZ experiment, the electroweak gauginos would need to have masses of at least $15$ TeV. Cases with very heavy gauginos have also been studied (see, e.g., [34]), as they provide a benchmark for inelastic direct-detection. However, such a heavy spectrum would lead to sizable loop-induced corrections to Eq. 1 if we assume the gluino to lie in the same mass range as the electroweakinos (see, e.g., discussion in [7]). In addition, the loop corrections to the Higgs mass would tend to become too large.
@@ V Sterile neutrino constraints
Figure 5: Mixing-angle–mass plane for a right-handed (sterile) neutrino: $\sin^{2}\!\bigl(2\theta\bigr)$ vs $m_{N}\,(\mathrm{GeV})$. The model points are distributed based on the total relic abundance by freeze-in and freeze-out. The background colour map shows the sterile-neutrino relic abundance $\Omega_{N}h^{2}$ from the Dodelson–Widrow mechanism.
@@
There are several direct and indirect constraints on right-handed (sterile) neutrinos. Collider searches [35, 36] and studies of kinematical end points in nuclear decays typically require large mixing between active and sterile neutrinos. Such a mixing cannot be realised within the type-I seesaw model and would require so large Yukawa couplings in the inverse and linear seesaw models that the sneutrino component would lead to overabundance in our scenario. In addition, short baseline oscillation experiments rule out very light sterile neutrinos below a few eV [37].
The indirect constraints are more severe. Light sterile neutrinos are relatively long-lived. They decay to active neutrinos via $N\rightarrow 3\nu$ if no other channels are kinematically available. The consequences of this decay depend on the sterile-neutrino lifetime. If the lifetime is of the order of one second, the decays will produce active neutrinos carrying energies higher than those expected from the thermal distribution, which can have an impact on Big Bang nucleosynthesis (BBN) [38]. Successful BBN therefore requires the sterile neutrinos either to decay well before BBN, allowing neutrinos to thermalise, or to be sufficiently long-lived.
If the sterile neutrinos have a lifetime longer than the age of the Universe, they constitute a (decaying) part of the dark matter. The dominant mechanism for sterile-neutrino dark matter generation is oscillations between active and sterile neutrinos known as the Dodelson–Widrow mechanism [19]. In this case, the relic density depends primarily on the sterile-neutrino mass and the mixing angle between the active and sterile flavours. The relic density can be approximately given by [39, 40]
@@
where $m_{N}$ is the sterile-neutrino mass and $\theta$ the mixing angle between the active and sterile neutrinos. The relic density depends rather mildly on model parameters since increasing the sterile-neutrino mass leads to a smaller mixing angle.
For all our benchmarks, we find that the lifetime of sterile neutrinos is of the order $10^{42}$ GeV-1 ($2.087\times 10^{10}$ yr), which is more than the age of the Universe ($1.38\times 10^{10}$ yr). Thus, they are stable on cosmological scales and contribute to the DM abundance.
As discussed earlier, the production of the right amount of sneutrino dark matter through freeze-in requires the neutrino Yukawa couplings of the order $\mathcal{O}(10^{-12})$. Eq. 5 implies that $m_{N}=(y^{\nu}v\sin\beta)^{2}/2m_{\nu}$. The heaviest active neutrino must be at least $0.04$ eV which leads to sterile-neutrino masses in the range 552020 eV, not much above the lowest allowed values by short baseline oscillation experiments. The associated Dirac mass term is of the order of one electronvolt which leads to an active-sterile mixing angle $\theta\gtrsim 10^{-2}$.
Linear and inverse seesaw models allow us to suppress this mixing by letting the sterile neutrinos be more massive. In the linear seesaw model, however, this comes with a price as this introduces additional interactions between the sterile sneutrinos and SM particles, which makes the freeze-in contribution too large as discussed earlier.
In the inverse seesaw model, the mass matrix has the form
@@
If we wish to have heavier sterile neutrinos while keeping the light neutrino mass constant, the factor $\mu_{S}/M_{N}$ must be greater than 1. The more we increase the sterile-neutrino mass, the larger $\mu_{S}$ must become. For proper suppression of the mixing $\mu_{S}/M_{N}\gtrsim 5$ would be required. When $\mu_{S}$ is significantly larger than $M_{N}$, the lower $2\times 2$ block of the inverse seesaw mass matrix becomes equivalent to that of the type-I seesaw model. Then we get sterile-neutrino masses $M_{N}^{2}/\mu_{S}$ for $N$ and $\mu_{S}$ for $S$, which means that we end up having the top $2\times 2$ block and type-I seesaw model. Hence, the inverse seesaw model too cannot successfully suppress the Dodelson–Widrow contribution to the relic density. The only model where there is no problem is that of Dirac neutrinos, since there are no right-handed neutrinos.
In Fig. 5, we plot the mass of the sterile neutrinos (with the largest mixing angle) against $\sin^{2}2\theta$ using points from the type-I seesaw model. The red dots correspond to points that yield the correct relic density with an error margin of $10\%$ accounting for numerical or package-related uncertainties. The black square and blue cross indicate benchmarks that lead to DM overabundance and underabundance, respectively. As the figure shows, all the points predict a relic density greater than 3, $\Omega_{N}h^{2}>3$, denoted by the reference line.
Another feature visible in Fig. 5 is the parabolic void in the lower left region of the plot, due to the lower bound on the Yukawa coupling specified in Table 1. By reducing this lower limit by half, this parabolic void disappears. Nonetheless, such points are excluded since they predict excessive contribution to DM relic density from sterile neutrinos via the Dodelson–Widrow mechanism.
The overabundance is mainly driven by the heaviest right-handed neutrino. For example, $m_{N}=50$ eV and $\sin^{2}2\theta\gtrsim 10^{-4}$ yields $\Omega h^{2}\gtrsim 2$. In the type-I seesaw, it is impossible to suppress this mixing while preserving a viable neutrino mass spectrum. Fig. 6 illustrates this through a numerical scan, where the Yukawa coupling values are now relaxed to $10^{-15}$ and the sterile-neutrino masses down to $0.1~\mathrm{eV}$. Points that allow smaller dark matter contributions from the right-handed neutrinos are excluded by oscillation experiments (lower right region), while the allowed points lead to a clear overabundance (upper middle). The experiments included in the blue contour are Bugey-3[41], NEOS[42], DANSS[43], Daya Bay[44], PROSPECT-I[45] and KATRIN[46].
In summary, although the combination of freeze-out and freeze-in can yield an acceptable relic density, the sterile-neutrino contribution is always too large. If sterile neutrinos were made sufficiently heavy with a sufficiently short lifetime, the couplings between the SM and sneutrinos would become too large to be viable.
@@ VI Conclusions
We showed that an acceptable relic density requires very small neutrino Yukawa couplings and a sterile-neutrino mass scale so low that the sterile neutrinos become cosmologically stable. In the region of parameter space where the relic density from the superpartners is viable, the production of sterile neutrinos in the early Universe via the Dodelson–Widrow mechanism leads to overabundance. Therefore, a scenario with freeze-in dark matter from sterile sneutrinos at the electroweak scale does not work for type-I, inverse, or linear seesaw mechanisms. For Dirac neutrinos, however, this scenario remains possible, as there are no sterile-neutrino mass eigenstates that would lead to overabundance.
@@ Acknowledgments
T.G. acknowledges Magnus Ehrnrooth Foundation for funding his doctoral studies. M.H. acknowledges the support from the Research Council of Finland (Grants #342777 and No.#371542). H.W. acknowledges the support from Carl Trygger Foundation (Grant CTS18:164) and Ruth and Nils-Erik Stenbäck Foundation. H.W. also thanks Helsinki Institute of Physics for hospitality during this project. The authors wish to thank the Finnish Computing Competence Infrastructure (FCCI) for supporting this project with computational and data storage resources.